2 mole of ideal monatomic gas at temperature T mixed with another 4 moles ideal monatomic gas at temperature 2 T . Then temperature of mixture

Grade:

1 Answers

neeraj agarwal
10 years ago
n Cv dT= n1 cv T1 + n2 cv T2 T` = n1T1 + n2T2 / n T`= 2T+8T / 6 T` = 5T/3